# Improved Stochastic Optimization of LogSumExp

### v1

@@ 3.2 Distributionally robust optimization with KL divergence
One of the approaches to training a model that is robust to data distribution shifts and noisy observations is called Distributionally Robust Optimization (DRO) (Kuhn et al., 2024). In contrast to the standard Empirical Risk Minimization (ERM) approach, which minimizes the average loss on the training sample, DRO minimizes the risk for the worst-case distribution among those close to a reference measure (e.g., empirical distribution). A prominent example is KL divergence DRO (Hu & Hong, ), which is formulated as the saddle-point problem
@@
However, when $n$ is large, computing the full gradient $\nabla\mathcal{L}(\theta)=\sum_{i=1}^{n}p_{i}^{*}(\theta)\nabla\ell_{i}(\theta)$ becomes costly. A straightforward approach (Levy et al., 2020) is to sample a batch $D$, compute the respective softmax weights $p_{i}^{D}(\theta)\coloneqq\frac{e^{\ell_{i}(\theta)/\lambda}}{\sum_{j\in D}e^{\ell_{j}(\theta)/\lambda}}$, and define a gradient estimator by
@@
$\min_{\begin{subarray}{c}\theta\in\Theta\\ \alpha\in\mathbb{R}\end{subarray}}\,G(\theta,\alpha)\coloneqq\frac{1}{n}\sum_{i=1}^{n}\Bigl\{\alpha+\frac{\lambda}{\rho}\log\bigl(1+\rho e^{(\ell_{i}(\theta)-\alpha)/\lambda}\bigr)\Bigr\}.$
@@
Consider California housing dataset (Pace & Barry, 1997) consisting of 20640 objects represented by 8 features. Let $\ell_{i}$ be the squared error of a linear model, $\ell_{i}(\theta)=(y_{i}-\theta^{\top}x_{i})^{2}$. We use SGD with Nesterov acceleration and the gradient estimator (18) (Levy et al., 2020) as the baseline approach for solving (17). As an alternative, accelerated SGD is used with our proposed gradient estimator (19). We considered values of the stepsize $\eta\in\{10^{-8},10^{-7},\ldots,10^{-4}\}$ and approximation accuracy parameter $\rho\in\{10^{-4},10^{-3},10^{-2}\}$. Momentum was set to 0.9 (without tuning) and entropy penalty $\lambda$ to 1. Least squares solution was used as a starting point for optimization.
(a) Batch size 10
(b) Batch size 100
(c) Batch size 1000
Figure 3: Convergence of SGD with Nesterov acceleration on the California housing dataset for the baseline gradient estimator (18) and the proposed estimator (19). The plots show $\mathcal{L}(\theta)$ versus the number of epochs for the best hyperparameters for different batch sizes. Error bars indicate $\pm$ one standard deviation over 20 runs.
Convergence curves for both approaches corresponding to the best hyperparameters (i.e., resulting in smallest objective values) are displayed in Figure 3. Y-axis depicts the value of $\mathcal{L}(\theta)$, X-axis displays the number of epochs (i.e., full passes over the dataset), error bars denote $\pm$ one standard deviation across 20 runs. We omit first few epochs to zoom in on smaller objective values. As seen from the figure, convergence of the proposed approach is consistently good across all batch sizes. As for the baseline, $|D|=10$ requires small stepsize $\eta=10^{-6}$ to avoid divergence, and thus results in slow convergence. Larger batches enable the increase of $\eta$ up to $10^{-5}$. The baseline still converges slower than the proposed approach with $|D|=100$, but almost matches its performance for $|D|=1000$.
@@ 3.3 Distributionally robust optimization with unbalanced OT
In the KL divergence DRO described in the previous subsection, uncertainty set is limited to distributions with the same support as the empirical measure $\mu=\frac{1}{n}\sum_{i}\delta_{x_{i}}$. Another popular approach, Wasserstein DRO (WDRO) (Mohajerin Esfahani & Kuhn, 2018; Sinha et al., 2020), considers the worst-case risk over shifts within a Wasserstein (OT) ball around a reference measure $\mu$ instead of the KL-ball in (16), thus including continuous probability measures. Unfortunately, this approach is not resilient to outliers that are geometrically far from the clean distribution since OT metric is sensitive to them (Nietert et al., 2023). A natural generalization is to switch to semi-balanced OT (Liero et al., 2018; Chizat et al., 2019; Kondratyev et al., 2016), which replaces a hard constraint on one of the marginals with a mismatch penalty function, e.g.,
$W_{\beta}(\nu,\mu)=\inf_{\begin{subarray}{c}\pi\in\mathcal{P}(\mathcal{X}\times\mathcal{X})\\ \pi_{1}=\nu\end{subarray}}\int c(x,z)\,\mathrm{d}\pi(x,z)+\beta\,D_{KL}(\pi_{2},\mu),$
@@
$\min_{\theta\in\Theta}\,\max_{\nu\in\mathcal{P}(\mathcal{X})}\;\int\ell(\theta,x)\,\mathrm{d}\nu(x)-\lambda W_{\beta}(\nu,\mu),$
where $\lambda>0$ is the Lagrangian penalty parameter. Using standard duality, Wang et al. (2024) showed that when $\mu=\frac{1}{n}\sum_{i}\delta_{x_{i}}$ is the empirical distribution, this is equivalent to
@@
where $\rho>0$ is a parameter controlling the accuracy of the approximation. This approximation can be efficiently optimized with SGD. Note that our method can also be applied to other DRO algorithms such as Sinkhorn DRO (Wang et al., 2021), which we omit to avoid redundancy.
@@
We consider MNIST dataset (Deng, 2012) with train and validation labels corrupted by feature-dependent noise (see Algan & Ulusoy, 2020) (noise ratio 25%), and original (clean) test labels. Let $\theta$ denote weights of a CNN with two convolutional layers (32 and 64 channels, kernel size 3, ReLU activations, and $\times2$ max pooling), followed by a fully connected classifier with one hidden layer of 128 units, and let $\ell(\theta;z)$ be its cross entropy loss on object $z$. In the experiment, SGD (with batch size 1) is applied to problems (21) (baseline) and (22) (proposed approach). We consider values of the stepsize $\eta\in\{10^{-6},10^{-5},10^{-4},10^{-3}\}$. Parameters $\lambda$ and $\beta$ were set to 1 since smaller values required a smaller stepsize and resulted in a slow convergence, while larger values pushed the model towards fitting the noisy distribution instead of the true one. Approximation accuracy parameter $\rho$ in (22) was set to 0.1. For the inner maximization problem in (20), just 5 iterations of Nesterov’s accelerated gradient method were sufficient to reach plateau in terms of the objective value. Additionally, we used SGD for the usual empirical risk minimization (ERM) to observe the effects of conventional (non-robust) training on noisy data.
@@ Appendix A Proofs for Section 2
$F_{\rho}(\varphi;\mu)=\inf_{\alpha\in\mathbb{R}}\alpha-1+\int e^{\varphi(x)-\alpha}g\left(\rho e^{\varphi(x)-\alpha}\right)\,\mathrm{d}\mu(x).$
@@
$F(\varphi;\mu)=\inf_{\alpha\in\mathbb{R}}\alpha-1+\int e^{\varphi(x)-\alpha}\,\mathrm{d}\mu(x).$
@@
$\int\ln\left(1+\rho e^{\varphi(x)-\alpha_{\rho}}\right)\,\mathrm{d}\mu(x)$ $=-\int\ln\left(1-\frac{\rho e^{\varphi(x)-\alpha_{\rho}}}{1+\rho e^{\varphi(x)-\alpha_{\rho}}}\right)\,\mathrm{d}\mu(x)$
$\geq-\ln\left(1-\int\frac{\rho e^{\varphi(x)-\alpha_{\rho}}}{1+\rho e^{\varphi(x)-\alpha_{\rho}}}\,\mathrm{d}\mu(x)\right)=-\ln(1-\rho),$
@@
$F_{\rho}(\varphi;\mu)=\alpha_{\rho}-1+\frac{1}{\rho}\int\ln\left(1+\rho e^{\varphi(x)-\alpha_{\rho}}\right)\,\mathrm{d}\mu(x)\geq\alpha_{\rho}-1-\frac{\ln(1-\rho)}{\rho}\geq\alpha_{\rho}+\frac{\rho}{2}.$ (23)
@@
$\int\frac{\rho e^{\varphi(x)-\alpha}}{1+\rho e^{\varphi(x)-\alpha}}\,\mathrm{d}\mu(x)\geq\rho.$
@@
$\int\frac{\rho e^{\varphi(x)-\alpha}}{1+\rho e^{\varphi(x)-\alpha}}\,\mathrm{d}\mu(x)\geq\int\left(\rho e^{\varphi(x)-\alpha}-\rho^{2}e^{2\varphi(x)-2\alpha}\right)\,\mathrm{d}\mu(x)=\rho e^{F(\varphi;\mu)-\alpha}-\rho^{2}e^{F(2\varphi;\mu)-2\alpha}.$
@@
$\int\ln\left(1+\rho e^{\varphi(x)-\alpha}\right)\,\mathrm{d}\mu(x)\geq\int e^{\varphi(x)-M}\ln\left(1+\rho e^{M-\alpha}\right)\,\mathrm{d}\mu(x)=e^{F(\varphi;\mu)-M}\ln\left(1+\rho e^{M-\alpha}\right)$
@@
$\operatorname{LogSumExp}(a_{1},\dots,a_{n})=\ln n+\ln\left(\int x\,\mathrm{d}\mu_{n}(x)\right)=\ln n+F(id;\mu_{n})$
@@
$\inf_{\alpha\in\mathbb{R}}\alpha-1+\frac{1}{\rho}\sum_{i=1}^{n}\ln(1+\rho e^{a_{i}-\alpha})$ $=\inf_{\alpha\in\mathbb{R}}\alpha-1+\frac{n}{\rho}\int\ln(1+\rho e^{x-\alpha})\,\mathrm{d}\mu_{n}(x)$
$=\inf_{\alpha\in\mathbb{R}}\alpha-1+\frac{n}{\rho}\int\ln\left(1+\frac{\rho}{n}e^{x-\alpha+\ln n}\right)\,\mathrm{d}\mu_{n}(x)$
@@
$\lambda F_{\rho}(\varphi/\lambda;\mu)$ $\geq\lambda\inf_{\alpha\in\mathbb{R}}\alpha-1+\frac{1}{\rho}\int\left(\ln\rho+\frac{\varphi(x)}{\lambda}-\alpha\right)_{+}\,\mathrm{d}\mu(x)$
$=\lambda(\ln\rho-1)+\inf_{\alpha\in\mathbb{R}}\alpha+\frac{1}{\rho}\int(\varphi(x)-\alpha)_{+}\,\mathrm{d}\mu(x).$
@@
Appendix B Additional details on the eOT experiment
B.1 SGD iterates for the dual and approximate semi-dual formulations
By the property of RKHS, if $u\in\mathcal{H}$, then $u(x)=\langle u,\kappa(\cdot,x)\rangle_{\mathcal{H}}$, and the derivatives of $f_{\varepsilon}$ are as follows:

### v2

@@ 3.2 Distributionally robust optimization with KL divergence
One of the approaches to training a model that is robust to data distribution shifts and noisy observations is called Distributionally Robust Optimization (DRO) (kuhnDistributionallyRobustOptimization2024). In contrast to the standard Empirical Risk Minimization (ERM) approach, which minimizes the average loss on the training sample, DRO minimizes the risk for the worst-case distribution among those close to a reference measure (e.g., empirical distribution). A prominent example is KL divergence DRO (huKullbackLeiblerDivergenceConstrained), which is formulated as the saddle-point problem
@@
However, when $n$ is large, computing the full gradient $\nabla\mathcal{L}(\theta)=\sum_{i=1}^{n}p_{i}^{*}(\theta)\nabla\ell_{i}(\theta)$ becomes costly. A straightforward approach (levy2020large) is to sample a batch $D$, compute the respective softmax weights $p_{i}^{D}(\theta)\coloneqq\frac{e^{\ell_{i}(\theta)/\lambda}}{\sum_{j\in D}e^{\ell_{j}(\theta)/\lambda}}$, and define a gradient estimator by
@@
$\min_{\begin{subarray}{c}\theta\in\Theta\\ \alpha\in\mathbb{R}\end{subarray}}\,G(\theta,\alpha)\coloneqq\frac{1}{n}\sum_{i=1}^{n}\Bigl\{\alpha+\frac{\lambda}{\rho}\log\bigl(1+\rho e^{(\ell_{i}(\theta)-\alpha)/\lambda}\bigr)\Bigr\}.$ (19)
@@
Consider the California housing dataset (pace1997sparse) consisting of 20,640 objects represented by 8 features. Let $\ell_{i}$ be the squared error of a linear model, $\ell_{i}(\theta)=(y_{i}-\theta^{\top}x_{i})^{2}$. We use accelerated SGD with the gradient estimator (18) (levy2020large) as the baseline approach for solving (17), and compare it to our proposed gradient estimator (20). We consider various values of the penalty coefficient $\lambda\in\{1/5,1,5\}$ and batch sizes $|D|\in\{10,10^{2},10^{3}\}$. For each configuration, we select the optimal learning rate from $\{10^{-9},10^{-8},\ldots,10^{-4}\}$. The approximation accuracy parameter $\rho$ in our method is varied across $\{10^{-1},10^{-3},10^{-5}\}$. Momentum is fixed at 0.9 (without tuning), and the least squares solution is used as the initial point for optimization.
Table 1: Objective value (17) (mean $\pm$ std across 10 runs) at epoch 50 for baseline (18) (levy2020large) and proposed gradient estimator (20) with different $\rho$ values. Results are shown for various penalty coefficients $\lambda$ and batch sizes $|D|$, with optimal learning rates selected from $\{10^{-9},\dots,10^{-4}\}$. Best results per column are shown in bold.
$\lambda=1/5$ $\lambda=1$ $\lambda=5$
Approach |D|=10|D|\!=\!10 $|D|\!=\!10^{2}$ $^{3} |D|=10|D|\!=\!10$ $|D|\!=\!10^{2}$ $^{3} |D|=10|D|\!=\!10$ $|D|\!=\!10^{2}$ $|D|\!=\!10^{3}$
Baseline (18) $26.9\!\pm\!0.7$ $15.6\!\pm\!6.0$ $\bm{9.1\!\pm\!4.6}$ $20.0\!\pm\!0.9$ $5.2\!\pm\!2.9$ $\bm{2.3\!\pm\!0.2}$ $7\!\pm\!0.01 0.88±0$$8\!\pm\!0.00$ $0.79\!\pm\!0.00$
(20), $\rho\!=\!10^{-1}$ $27.7\!\pm\!0.6$ $27.7\!\pm\!0.7$ $40.1\!\pm\!0.5$ $21.1\!\pm\!1.1$ $21.3\!\pm\!1.1$ $21.8\!\pm\!2.1$ $0.87\!\pm\!0.01$ $7\!\pm\!0.01 0.88±0$$8\!\pm\!0.02$
(20), $\rho\!=\!10^{-3}$ $21.2\!\pm\!9.8$ $18.6\!\pm\!7.7$ $25.3\!\pm\!0.1$ $\bm{2.1\!\pm\!0.0}$ $\bm{2.1\!\pm\!0.0}$ $\bm{2.5\!\pm\!1.2}$ $\bm{0.76\!\pm\!0.02}$ $\bm{0.78\!\pm\!0.00}$ $\bm{0.78\!\pm\!0.00}$
(20), $\rho\!=\!10^{-5}$ $19.2\!\pm\!9.6$ $17.5\!\pm\!6.6$ $24.3\!\pm\!0.3$ $3.0\!\pm\!0.0$ $3.0\!\pm\!0.0$ $3.0\!\pm\!0.0$ $1.03\!\pm\!0.00$ $1.03\!\pm\!0.00$ $1.03\!\pm\!0.00$
Numerical results are presented in Table 1, showing the objective value (mean $\pm$ standard deviation across 10 runs) after 50 epochs, where the methods typically reach a plateau. In each column, the best-performing configurations are highlighted in bold. For $\lambda=1/5,|D|\in\{10,10^{2}\}$, no results are displayed in bold as all configurations perform similarly. As seen from the table, the baseline and our estimator achieve comparable performance for large batch sizes ($|D|=10^{3}$). However, for smaller batches, our method typically outperforms the baseline. Both approaches handle various $\lambda$ values well, with the exception of the baseline method combined with small batch sizes.
Regarding the approximation parameter $\rho$, large values ($\rho=10^{-1}$) generally result in a noticeable approximation gap, while excessively small values ($\rho=10^{-5}$) deteriorate the smoothness of the objective and consequently slow convergence. The intermediate value $\rho=10^{-3}$ thus provides the best trade-off in this experiment, offering both good approximation accuracy and favorable optimization properties.
@@ 3.3 Distributionally robust optimization with unbalanced OT
In the KL divergence DRO described in the previous subsection, uncertainty set is limited to distributions with the same support as the empirical measure $\mu=\frac{1}{n}\sum_{i}\delta_{x_{i}}$. Another popular approach, Wasserstein DRO (WDRO) (mohajerinesfahaniDatadrivenDistributionallyRobust2018; sinhaCertifyingDistributionalRobustness2020), considers the worst-case risk over shifts within a Wasserstein (OT) ball around a reference measure $\mu$ instead of the KL-ball in (16), thus including continuous probability measures. Unfortunately, this approach is not resilient to outliers that are geometrically far from the clean distribution since OT metric is sensitive to them (nietert2023outlier). A natural generalization is to switch to semi-balanced OT (liero_optimal_2018; chizat_unbalanced_2019; kondratyevNewOptimalTransport2016), which replaces a hard constraint on one of the marginals with a mismatch penalty function, e.g.,
$W_{\beta}(\nu,\mu)=\inf_{\begin{subarray}{c}\pi\in\mathcal{P}(\mathcal{X}\times\mathcal{X})\\ \pi_{1}=\nu\end{subarray}}\int_{{\mathcal{X}\times\mathcal{X}}}c(x,z)\,\mathrm{d}\pi(x,z)+\beta\,D_{KL}(\pi_{2},\mu),$
@@
$\min_{\theta\in\Theta}\,\max_{\nu\in\mathcal{P}(\mathcal{X})}\;\int_{{\mathcal{X}}}\ell(\theta,x)\,\mathrm{d}\nu(x)-\lambda W_{\beta}(\nu,\mu),$
where $\lambda>0$ is the Lagrangian penalty parameter. Using standard duality, wang2024outlier showed that when $\mu=\frac{1}{n}\sum_{i}\delta_{x_{i}}$ is the empirical distribution, this is equivalent to
@@
where $\rho>0$ is a parameter controlling the accuracy of the approximation. This approximation can be efficiently optimized with SGD. Note that our method can also be applied to other DRO algorithms such as Sinkhorn DRO (wang2021sinkhorn), which we omit to avoid redundancy.
@@
We consider MNIST dataset (deng2012mnist) with train and validation labels corrupted by feature-dependent noise (see algan2020label) (noise ratio 25%), and original (clean) test labels. Let $\theta$ denote weights of a CNN with two convolutional layers (32 and 64 channels, kernel size 3, ReLU activations, and $\times2$ max pooling), followed by a fully connected classifier with one hidden layer of 128 units, and let $\ell(\theta;z)$ be its cross entropy loss on object $z$. In the experiment, SGD (with batch size 1) is applied to problems (22) (baseline) and (23) (proposed approach). We consider values of the stepsize $\eta\in\{10^{-6},10^{-5},10^{-4},10^{-3}\}$. Parameters $\lambda$ and $\beta$ were set to 1 since smaller values required a smaller stepsize and resulted in a slow convergence, while larger values pushed the model towards fitting the noisy distribution instead of the true one. Approximation accuracy parameter $\rho$ in (23) was set to 0.1. For the inner maximization problem in (21), just 5 iterations of Nesterov’s accelerated gradient method were sufficient to reach plateau in terms of the objective value. Additionally, we used SGD for the usual empirical risk minimization (ERM) to observe the effects of conventional (non-robust) training on noisy data.
@@ Appendix A Proofs for Section 2
$F_{\rho}(\varphi;\mu)=\inf_{\alpha\in\mathbb{R}}\alpha-1+\int_{{\mathcal{X}}}e^{\varphi(x)-\alpha}g\left(\rho e^{\varphi(x)-\alpha}\right)\,\mathrm{d}\mu(x).$
@@
$F(\varphi;\mu)=\inf_{\alpha\in\mathbb{R}}\alpha-1+\int_{{\mathcal{X}}}e^{\varphi(x)-\alpha}\,\mathrm{d}\mu(x).$
@@
$\int_{{\mathcal{X}}}\ln\left(1+\rho e^{\varphi(x)-\alpha_{\rho}}\right)\,\mathrm{d}\mu(x)$ $=-\int_{{\mathcal{X}}}\ln\left(1-\frac{\rho e^{\varphi(x)-\alpha_{\rho}}}{1+\rho e^{\varphi(x)-\alpha_{\rho}}}\right)\,\mathrm{d}\mu(x)$
$\geq-\ln\left(1-\int_{{\mathcal{X}}}\frac{\rho e^{\varphi(x)-\alpha_{\rho}}}{1+\rho e^{\varphi(x)-\alpha_{\rho}}}\,\mathrm{d}\mu(x)\right)=-\ln(1-\rho),$
@@
$F_{\rho}(\varphi;\mu)=\alpha_{\rho}-1+\frac{1}{\rho}\int_{{\mathcal{X}}}\ln\left(1+\rho e^{\varphi(x)-\alpha_{\rho}}\right)\,\mathrm{d}\mu(x)\geq\alpha_{\rho}-1-\frac{\ln(1-\rho)}{\rho}\geq\alpha_{\rho}+\frac{\rho}{2}.$ (24)
@@
$\int_{{\mathcal{X}}}\frac{\rho e^{\varphi(x)-\alpha}}{1+\rho e^{\varphi(x)-\alpha}}\,\mathrm{d}\mu(x)\geq\rho.$
@@
$\int_{{\mathcal{X}}}\frac{\rho e^{\varphi(x)-\alpha}}{1+\rho e^{\varphi(x)-\alpha}}\,\mathrm{d}\mu(x)\geq\int_{{\mathcal{X}}}\left(\rho e^{\varphi(x)-\alpha}-\rho^{2}e^{2\varphi(x)-2\alpha}\right)\,\mathrm{d}\mu(x)=\rho e^{F(\varphi;\mu)-\alpha}-\rho^{2}e^{F(2\varphi;\mu)-2\alpha}.$
@@
$\int_{{\mathcal{X}}}\ln\left(1+\rho e^{\varphi(x)-\alpha}\right)\,\mathrm{d}\mu(x)\geq\int_{{\mathcal{X}}}e^{\varphi(x)-M}\ln\left(1+\rho e^{M-\alpha}\right)\,\mathrm{d}\mu(x)=e^{F(\varphi;\mu)-M}\ln\left(1+\rho e^{M-\alpha}\right)$
@@
$\operatorname{LogSumExp}(a_{1},\dots,a_{n})=\ln n+\ln\left(\int_{{\mathcal{X}}}x\,\mathrm{d}\mu_{n}(x)\right)=\ln n+F(id;\mu_{n})$
@@
$\inf_{\alpha\in\mathbb{R}}\alpha-1+\frac{1}{\rho}\sum_{i=1}^{n}\ln(1+\rho e^{a_{i}-\alpha})$ $=\inf_{\alpha\in\mathbb{R}}\alpha-1+\frac{n}{\rho}\int_{{\mathcal{X}}}\ln(1+\rho e^{x-\alpha})\,\mathrm{d}\mu_{n}(x)$
$=\inf_{\alpha\in\mathbb{R}}\alpha-1+\frac{n}{\rho}\int_{{\mathcal{X}}}\ln\left(1+\frac{\rho}{n}e^{x-\alpha+\ln n}\right)\,\mathrm{d}\mu_{n}(x)$
@@
$\lambda F_{\rho}(\varphi/\lambda;\mu)$ $\geq\lambda\inf_{\alpha\in\mathbb{R}}\alpha-1+\frac{1}{\rho}\int_{{\mathcal{X}}}\left(\ln\rho+\frac{\varphi(x)}{\lambda}-\alpha\right)_{+}\,\mathrm{d}\mu(x)$
$=\lambda(\ln\rho-1)+\inf_{\alpha\in\mathbb{R}}\alpha+\frac{1}{\rho}\int_{{\mathcal{X}}}(\varphi(x)-\alpha)_{+}\,\mathrm{d}\mu(x).$
@@
Proof of Lemma 2.7.
Recall that we have
$f_{\rho}(t)=\frac{1}{\rho}\!\left((\rho t)\log(\rho t)+(1-\rho t)\log(1-\rho t)\right)+1-t\log\rho.$
Simplifying, we obtain
$f_{\rho}(t)=t\log t+\frac{1}{\rho}(1-\rho t)\log(1-\rho t)+1.$
The first derivative is calculated as follows:
$\frac{d}{dt}\big(t\log t\big)=\log t+1,\qquad\frac{d}{dt}\!\left(\frac{1}{\rho}(1-\rho t)\log(1-\rho t)\right)=-\big(\log(1-\rho t)+1\big),$
so
$f_{\rho}^{\prime}(t)=(\log t+1)-\big(\log(1-\rho t)+1\big)=\log t-\log(1-\rho t)=\log\!\left(\frac{t}{1-\rho t}\right).$
The second derivative is calculated as follows:
$\frac{d}{dt}\!\left(\log t\right)=\frac{1}{t},\qquad\frac{d}{dt}\!\left(\log(1-\rho t)\right)=-\frac{\rho}{1-\rho t},$
thus
$f_{\rho}^{\prime\prime}(t)=\frac{1}{t}+\frac{\rho}{1-\rho t}=\frac{1}{t(1-\rho t)}.$
By symmetry, we can see that the minimum value of the second derivative is achieved at $t^{*}=\frac{1}{2\rho}$, and it is equal to $4\rho$. Thus, for all $t\in{\rm dom}f_{\rho}$, we have that $f_{\rho}^{\prime\prime}(t)\geq 4\rho>\rho$. Thus, by (nesterov2018lectures, Theorem 2.1.11), $f_{\rho}$ is $\rho$-strongly convex. By (Zhou2018FenchelDuality, Theorem 1) this also implies that its conjugate function $f_{\rho}^{*}$ is $\frac{1}{\rho}$-smooth. ∎
Appendix B Additional materials on entropic OT
B.1 Related works on eOT
This subsection provides an overview of selected works on continuous entropy-regularized optimal transport. In (genevay2016stochastic), the authors tackled this problem by introducing an RKHS and optimizing the dual function (12) with SGD. This approach was extended by seguy2018large, who parameterized the dual potentials with neural networks instead of an RKHS to improve scalability. Subsequently, daniels2021score leverage this approach to approximate the optimal transport plan, using it to develop a score-based generative model. Although this direction mostly results in computationally efficient methods that works with a general cost function, a key drawback is that small values of the regularization coefficient $\varepsilon$ cause numerical instabilities due to the exponential term in the dual objective; see Remark 3.1. The work by (korotin2023neural) studies a more general formulation known as weak OT. The authors formulate it as a maximin problem and develop a neural-network-based algorithm under the assumption of a quadratic cost, a restriction that is later relaxed in (asadulaev2024neural). However, these methods are computationally intensive due to their adversarial training nature. The paper by mokrov2024energyguided approaches eOT from the perspective of energy-based models. Unfortunately, the resulting solver is computationally expensive as it involves iterative Langevin dynamics . Another popular approach to eOT in recent years is via the Schrödinger bridge (SB), e.g., (gushchin2023entropic). While SB-based solvers are also often computationally intensive, a more cost-efficient solution has been proposed by (korotin2024light). However, it relies on the quadratic cost assumption and does not support general cost. We would also like to note that a promising direction for future work is leveraging our approach for minimizing the objective (8) in (korotin2024light) to further improve scalability.
B.2 Experiment with RKHS representation of dual potentials
As mentioned earlier, LSOT (seguy2018large) is inspired by the continuous eOT approach of genevay2016stochastic. This work considers a reproducing kernel Hilbert space (RKHS) $\mathcal{H}$ defined on $\mathcal{X}$, with a kernel $\kappa$, and applies SGD to solve the dual problem. Such approach suffers from the same numerical instability as LSOT, see Remark 3.1. As an alternative, we again consider the approximation (15) of the semi-dual objective which can also be maximized by SGD. Although the variable $\alpha$ is, in general, a function of $x$, we empirically found that tuning a common scalar value $\alpha\in\mathbb{R}$ for all samples works well in the experiments described below.
Analytic form of SGD iterates for both objectives can be derived as follows. By the property of RKHS, if $u\in\mathcal{H}$, then $u(x)=\langle u,\kappa(\cdot,x)\rangle_{\mathcal{H}}$. Therefore, the derivatives of $f_{\varepsilon}$ take the form